# Primordial Black Hole Reformation in the Early Universe

### v1

@@ Abstract
Primordial black holes (PBH) can arise in a wide range of scenarios, from inflation to first-order phase transitions. Light PBHs, such as those produced during preheating or at the GUT scale, could induce an early matter-dominated phase given a moderate initial abundance. During the early matter-domination, the growth of initial PBH density perturbations can trigger collapse on horizon scales, producing much heavier PBHs. While the remaining original PBHs evaporate and reheat the Universe, these massive reformed PBHs survive for an extended period of time, producing potentially observable signatures at the present. We study this PBH reformation scenario and show that those reformed PBHs can emit significant quantities of gamma rays detectable by the next generation of experiments. The rapid reheating after matter domination generates a coincident stochastic gravitational wave background, which could be within range of the upcoming CMB-S4 experiment. The PBH reformation scenario provides an intriguing possibility of decoupling the current PBH population and the initial formation mechanism from early Universe physics, while providing opportunities for observation through multi-messenger astronomy.
@@ I Introduction
Primordial black holes (PBHs) can be produced in a variety of new physics scenarios, ranging from the collapse of perturbations seeded by inflationary fluctuations Carr and Hawking (1974); Carr (1975); Garcia-Bellido et al. (1996); Yokoyama (1997); Garcia-Bellido and Ruiz Morales (2017); Musco et al. (2021) to first-order phase transitions Hawking et al. (1982); Moss (1994); Jung and Okui (2021); Hong et al. (2020); Kawana and Xie (2022); Lu et al. (2022); Kawana et al. (2022); Lu et al. (2023a, b); Marfatia et al. (2024); Liu et al. (2022); Kawana et al. (2023) and many others Hawking (1989); Cotner et al. (2018); Conzinu et al. (2020); Helfer et al. (2017); Ruffini and Bonazzola (1969). The conceivable mass range of PBHs span more than 40 orders of magnitude, from the Planck mass to supermassive black holes. Of particular interest are light PBHs with masses between $5\times 10^{14}g<M_{\rm PBH}\lesssim 10^{17}g$, which can emit copious amounts of high energy Hawking radiation Hawking (1974, 1976); Page (1976). Consequently, this parameter space has been constrained by a variety of experiments Carr et al. (2010, 2016); Boudaud and Cirelli (2019); Huang and Zhou (2024); Kim (2021); Laha et al. (2021) searching for this emission. The next generation of observatories Addazi et al. (2022); Acharya et al. (2018); Albert et al. (2019) will further improve the sensitivity to high energy $\gamma$ and cosmic rays, enabling them detect a small but significant population of evaporating PBHs.
@@
PBHs with masses below $5\times 10^{14}g$ are usually considered to have completely evaporated by the present day Page (1976); MacGibbon and Webber (1990) and could be detectable by their traces on the Cosmic Microwave Background (CMB) and nuclide abundances from Big Bang Nucleosynthesis (BBN) Carr et al. (2010); Acharya and Khatri (2020); Chluba et al. (2020). These bounds extend down to $M\gtrsim 10^{9}g$, below which the PBH lifetime is short enough to evaporate before the epoch of BBN $\sim 1s$. Recently, a series of papers Inomata et al. (2020); Papanikolaou et al. (2021); Domènech et al. (2021a, b) have shown that extremely light PBHs could usher in an era of early matter-domination (eMD) and source a large gravitational wave (GW) background when they evaporate, during the rapid transition from matter-domination to radiation-domination (RD). Although generally undetectable by present-day GW experiments, these GWs would constitute an excess contribution to the radiation energy density. As a result, $\Delta N_{\rm eff}$ limits derived from CMB and BBN observations constrain the abundance of PBHs down to $M\gtrsim 10^{-2}g$ Domènech et al. (2021a, b).
@@
In this paper, we consider PBH reformation during and after the eMD era. Shortly after matter-radiation equality, the overdensities associated with the PBHs grow and could become substantial before they evaporate. During the eMD phase, the probability of PBH reformation is enhanced, following a power law Khlopov and Polnarev (1980); Polnarev and Khlopov (1981); Harada et al. (2016, 2017); Kokubu et al. (2018) instead of the usual exponential dependence Carr (1975); Young et al. (2014). This enables the possibility that collections of PBHs may collapse and form much larger PBHs. A similar concept was explored in Ref. De Luca et al. (2023), where they discussed mergers of solar mass PBHs into supermassive black hole seeds, enhanced by strong clustering. In contrast, we focus on very light PBHs that would otherwise evaporate, and could initiate an eMD era in the early Universe, with collapse of a PBH gas rather than individual mergers as the formation mechanism. These reformed PBHs would be much lighter, actively emitting Hawking radiation which could be detected along with the coincident GW signal.
@@
We start by reviewing the relevant quantities and scales for PBH formation and the onset of an eMD era in Sec. II. The power spectrum and its growth is discussed in Sec. III which leads to the PBH reformation during eMD in Sec. IV. The accompanying GW background is discussed in Sec. V. Then in Sec. VI, we analyze PBH reformation after the reheating, introducing clustering effects during the initial formation. Finally, we conclude in Sec. VII.
@@ II PBH Parameters
$M_{\rm PBH}=\gamma\frac{4\pi M_{P}^{2}}{H}\simeq 47\leavevmode\nobreak\ g\leavevmode\nobreak\ \left(\frac{\gamma}{0.5}\right)\left(\frac{T}{10^{15}\textrm{ GeV}}\right)^{-2}\leavevmode\nobreak\ ,$ (1)
where we take $\gamma=0.5$ as a fiducial value Carr and Kuhnel (2020) for both the initial formation and the reformation. Unless otherwise specified, this relation is used for the following calculations, but should be modified if the initial formation scenario involves subhorizon scales (e.g. from bubble collisions Hawking et al. (1982); Moss (1994); Jung and Okui (2021) or compact remnants Hong et al. (2020); Kawana and Xie (2022); Kawana et al. (2022); Lu et al. (2022)).
At their formation, the fraction of PBHs with respect to the overall energy density is characterized by $\beta=\rho_{\rm PBH}/\rho_{\rm tot}$, where $\rho_{\rm tot}$ is the total energy density of the PBHs and the relativistic SM plasma. Here we differentiate between the density fraction of PBHs at initial formation, $\beta_{\rm if}$, and the density fraction at reformation $\beta_{\rm rf}$. We focus on initial formation masses, $M_{\rm pl}<M_{\rm if}\lesssim 4\times 10^{8}g$, which are unconstrained by BBN and CMB considerations. Such light PBHs are typical of GUT-scale formation Anantua et al. (2009); Zagorac et al. (2019) and preheating scenarios Bassett and Tsujikawa (2001); Green and Malik (2001); Martin et al. (2020).
Since PBHs are non-relativistic objects, their density increases relative to the SM plasma with redshift, which can result in an early matter-domination (eMD) phase. The matter(PBH)-radiation equality scale factor is simply related to the scale factor at initial formation by $a_{\rm eq}=a_{\rm if}/\beta_{\rm if}$. Likewise, the Hubble rate at the equality is proportional to the Hubble rate at initial formation, and the comoving horizon wavenumber is given by
$k_{\rm eq}=a_{\rm eq}H_{\rm eq}=\sqrt{2}a_{\rm if}H_{\rm if}\beta_{\rm if}\leavevmode\nobreak\ .$ (2)
@@
$\tau_{\rm evap}=4.0\times 10^{-4}s\left(\frac{M_{\rm PBH}}{10^{8}g}\right)^{3}\left(\frac{108}{g_{H}}\right)\leavevmode\nobreak\ ,$ (3)
where the Hawking radiation degrees of freedom $g_{H}\simeq 108$ incorporates all the SM modes for $M_{\rm PBH}\lesssim 10^{13}g$.
@@
$\beta_{\rm if}>4\times 10^{-6}\left(\frac{M_{\rm PBH}}{1g}\right)^{-1}\left(\frac{g_{*}(T_{\rm RH})}{g_{*}(T_{\rm if})}\right)^{1/3}\leavevmode\nobreak\ ,$ (4)
@@
×$\times\left(\frac{M_{\rm PBH}}{1\leavevmode\nobreak\ g}\right)^{-3/2}\left(\frac{g_{*}(T_{\rm RH})}{106.75}\right)^{-1/4}\left(\frac{g_{H}}{108}\right).$
@@
$=$ $2\times 10^{-4}\,a_{\rm if}H_{\rm if}\left(\frac{\beta}{\gamma}\right)^{1/3}\left(\frac{M_{\rm PBH}}{1\leavevmode\nobreak\ g}\right)^{-2/3}.$
@@ III Power Spectrum
The mean separation between PBHs at the initial formation, $\bar{r}$, defines the comoving ultraviolet cut-off wavenumber Papanikolaou et al. (2021); Domènech et al. (2021a),
$k_{\rm UV}=\frac{a}{\bar{r}}=\left(\frac{\beta_{\rm if}}{\gamma}\right)^{1/3}a_{\rm if}H_{\rm if}\leavevmode\nobreak\ .$ (7)
For $k<k_{\rm UV}$, the PBHs can be treated as a Poissonian fluid with initial density power spectrum Papanikolaou et al. (2021)
$\mathcal{P}_{\delta}(t_{\rm if})=\frac{2}{3\pi}\left(\frac{k}{k_{\rm UV}}\right)^{3}\leavevmode\nobreak\ .$ (8)
During the eMD, PBH density perturbations grow linearly with scale factor Meszaros (1974) once they cross the horizon. On the other hand, the growth of short wavelength modes that enter during RD is delayed until matter-radiation equality.
The evolution of the power spectrum is described by the transfer function, $\mathcal{P}_{\delta}(t)=\mathcal{T}^{2}(t)\mathcal{P}_{\delta}(t_{\rm if})$. We numerically computed the linear transfer function in Appendix A, and found a semi-analytic fit to the transfer function at the reheating time,
@@
$\simeq\begin{cases}\frac{3}{2}\,y,&\text{for $\kappa\gg 1$}\\ \frac{4}{15}\,\kappa^{2}y,&\text{for $\kappa\ll 1$}\end{cases}\leavevmode\nobreak\ .$
Here $\kappa=k/k_{\rm eq}$ and $y=a/a_{\rm eq}$ are the wavenumber and scale factor with respect to the matter-radiation equality, and $y\gg 1$ is assumed. We truncated the growth of the power spectrum at the density of a black hole, imposing the constraint
$\mathcal{P}_{\delta}(k)\lesssim\left(\frac{k}{aH}\right)^{4}\leavevmode\nobreak\ .$ (10)
However, even in the case of initial PBH clustering, we find that the power spectrum is well below this limit.
@@
$\sigma^{2}_{\delta}(r,t)=\int^{\infty}_{0}\mathcal{P}_{\delta}(t)W^{2}(kr)\frac{dk}{k}\leavevmode\nobreak\ .$ (11)
@@
$W^{2}(kr\gg 1)\simeq\frac{9}{2}\frac{1}{(kr)^{4}}\simeq\frac{9}{8}\frac{1}{\kappa^{4}y^{2}}\leavevmode\nobreak$ (13)
@@
$P_{\delta}(t_{\rm RH})W^{2}(kr)\propto\begin{cases}\kappa^{-1},&\text{for $\kappa>1$}\\ \kappa^{3},&\text{for $\kappa<1$}\end{cases}\leavevmode\nobreak$ (14)
with $y_{\rm RH}$ dependence canceled for a long enough period of eMD.
In the rest of the paper, we use this linear computation and leave non-linear studies to future works. Although there are well-known methods Peebles (1980); Hamilton et al. (1991); Mo et al. (1995); Peacock and Dodds (1996); Smith et al. (2003); Takahashi et al. (2012); Bartlett et al. (2024) to analytically transform the linear power spectrum into the non-linear power spectrum, we find that they do not yield reasonable results in this application. In general, we expect that non-linearities will only enhance the overall variance so our linear calculation could be considered a conservative estimate.
Figure 1: The initial PBH power spectrum (gray), power spectrum at the moment of reheating (black), and the variance at the horizon scale of reheating per logarithmic interval of wavenumber (red). The initial power spectrum is given by the initial Poisson distribution of PBHs, which then grows according to the transfer function, and then multiplied by the window function to give the variance. The parameters are chosen to produce a sizeable population of reformed PBHs, represented by Case A in Table 1 and Fig. 2. All three lines are truncated at the UV cutoff $k_{\rm UV}$ in Eq. (7).
In Fig. 1, we show the power spectrum at the initial formation time and at the time of reheating. As in Eq. (14), the variance per logarithmic wavenumber intervals levels off at the matter-radiation equality scale $k\simeq k_{\rm eq}$. If any initial clustering of PBHs exists, the power at small scales increases and the contribution to the variance continues to increase for larger wavenumbers, making the behavior around the ultraviolet cutoff dominate the reformation. We consider such a case in Sec. VI.
@@ IV PBH Reformation in eMD
Refs. Khlopov and Polnarev (1980); Polnarev and Khlopov (1981); Harada et al. (2016, 2017); Kokubu et al. (2018) considered the possibility of PBH formation during MD. In the absence of pressure forces, the critical overdensity threshold may be expected to tend toward zero Musco and Miller (2013). However, two additional considerations of isotropy and homogeneity are required to ensure gravitational collapse. Instead of forming a PBH, an anisotropic perturbation undergoes a “pancake” collapse, forming caustics and resulting in a virialized configuration. Using the anisotropy distribution of Ref. Doroshkevich (1970), the fraction of perturbations that are sufficiently spherically symmetric enough to collapse into a PBH was computed by Ref. Khlopov and Polnarev (1980); Polnarev and Khlopov (1981); Harada et al. (2016), well approximated by $\beta=0.05556\sigma_{\delta}^{5}$.
Additionally, these perturbations must be homogeneous enough to form a horizon (i.e. satisfy the hoop criterion) before forming a “naked singularity” which signifies the breakdown of the calculation. This limitation on the structure is satisfied by only a fraction of perturbations $\beta=3.70\sigma_{\delta}^{3/2}$ Kokubu et al. (2018), but may be model dependent Harada et al. (2016). Taken independently, these two conditions would suggest a minimum PBH formation probability proportional to $\beta\propto\sigma^{13/2}$. However, both conditions are more easily satisfied by larger perturbations, which can shrink to their larger Schwarzchild radii before becoming too anisotropic or inhomogenous. Since these conditions are correlated and dependent, the actual conditional probability should then be much higher than the minimum estimate. Therefore, we use the optimistic estimate of $\beta\simeq 0.05556\sigma^{5}$ in this paper, which assumes that perturbations satisfying the much more stringent constraint of isotropy are also homogeneous enough. We ignore spin and angular momentum effects Harada et al. (2017) which may alter the probability of PBH formation for small values of $\sigma_{\delta}$.
@@
Figure 2: Two populations of PBHs reformed during the eMD era that could be detected in the next generation of experiments. Both Case A (red) and Case B (blue) are in the allowed region of initial PBH parameters (green region) where the initial density $\beta_{\rm if}$ is large enough to result in an eMD phase and below the GW bound. The PBH spectrum for Case A lies below the current bounds from BBN and the emission of PBHs reformed in Case B could be detected by the next generation of CMB and $\gamma$-ray experiments. The gray-shaded regions are constraints from BBN Carr et al. (2010), CMB Acharya and Khatri (2020); Chluba et al. (2020), extragalactic $\gamma$-rays Carr et al. (2010), galactic $\gamma$-rays Carr et al. (2016), Voyager-1 $e^{\pm}$ Boudaud and Cirelli (2019), and AMS-02 Huang and Zhou (2024).
We apply the formula of Ref. Harada et al. (2016) to PBH reformation during the eMD, from matter-radiation equality up until evaporation and reheating. Evaluating the variance at the horizon size of each epoch by Eqs. (11) and (14), we find it to be roughly constant and approximated by
@@
$f_{\rm PBH}=4.0\times 10^{20}\beta\left(\frac{M_{\rm if}}{1\leavevmode\nobreak\ g}\right)^{-3/2}\left(\frac{g_{*}(T_{\rm RH})}{106.75}\right)^{1/12}\leavevmode\nobreak\ .$ (16)
This flat mass spectrum has endpoints defined by the horizon scale at matter-radiation equality and reheating, $k_{\rm RH}<k<k_{\rm eq}$. From Eqs. (1), (2), and (5), the mass range of the reformed PBH spectrum is
$\frac{M_{\rm if}}{\sqrt{2}\beta_{\rm if}^{2}}\leq M_{\rm RH}\leq 6.0\times 10^{10}g\left(\frac{\gamma}{0.5}\right)\left(\frac{M_{\rm if}}{1\leavevmode\nobreak\ g}\right)^{3}\leavevmode\nobreak\ .$ (17)
@@
For Case A, the reformed PBH population are at the cusp of current BBN constraints from observations of the primordial deuterium abundance Zel’dovich et al. (1977); Lindley (1980); Ellis et al. (1992). Experiments targeting the processes relevant for deuterium production could further reduce prediction uncertainties and lead to either more stringent constraints or an anomaly Mossa et al. (2020); Xu et al. (2023); Shen and He (2024). The reformed PBH population of Case B have lifetimes comparable to that of the Universe $\sim 10^{10}$ yr, and could be emitting high energy particles ($\gamma$,$e^{\pm}$) at the present. The evaporation signal of these reformed PBH may be within the sensitivity limit of the operational run of LHAASO Addazi et al. (2022); Yang et al. (2024) and next generation $\gamma$-ray experiments such as CTA Acharya et al. (2018) and SGSO Albert et al. (2019). We consider a possible coincident GW signature in the next section.

### v2

@@ Abstract
Primordial black holes (PBH) can arise in a wide range of scenarios, from inflation to first-order phase transitions. Light PBHs, such as those produced during preheating, in bounce cosmologies, or at the GUT scale, could induce an early matter-dominated phase given a moderate initial abundance. During the early matter domination, the growth of initial PBH density perturbations can trigger collapse on horizon scales, producing much heavier PBHs. While the remaining original PBHs evaporate and reheat the Universe, these massive reformed PBHs survive for an extended period of time, producing potentially observable signatures at the present. We study this PBH reformation scenario and show that those reformed PBHs can emit significant quantities of gamma rays detectable by the next generation of experiments. The rapid reheating after matter domination generates a coincident stochastic gravitational wave background, which could be within the range of the upcoming CMB-S4 experiment. The PBH reformation scenario provides an intriguing possibility of decoupling the current PBH population and the initial formation mechanism from early Universe physics, while providing opportunities for observation through multi-messenger astronomy.
@@ I Introduction
Primordial black holes (PBHs) can be produced in a variety of new physics scenarios, ranging from the collapse of perturbations seeded by inflationary fluctuations [1, 2, 3, 4, 5, 6] to first-order phase transitions [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and many others [20, 21, 22, 23, 24]. The conceivable mass range of PBHs span more than 40 orders of magnitude, from the Planck mass to supermassive black holes. Of particular interest are light PBHs with masses between $5\times 10^{14}\textrm{g}<M_{\rm PBH}\lesssim 10^{17}\textrm{g}$, which can emit copious amounts of high energy Hawking radiation [25, 26, 27]. Consequently, this parameter space has been constrained by a variety of experiments [28, 29, 30, 31, 32, 33] searching for this emission. The next generation of observatories [34, 35, 36] will further improve the sensitivity to high energy $\gamma$ and cosmic rays, enabling them detect a small but significant population of evaporating PBHs.
@@
PBHs with masses below $5\times 10^{14}\textrm{g}$ are usually considered to have completely evaporated by the present day [27, 37] and could be detectable by their traces on the Cosmic Microwave Background (CMB) and nuclide abundances from Big Bang Nucleosynthesis (BBN) [28, 38, 39]. These bounds extend down to $M\gtrsim 10^{9}\textrm{g}$, below which the PBH lifetime is short enough to evaporate before the epoch of BBN $\sim 1\textrm{s}$. Recently, a series of papers [40, 41, 42, 43, 44, 45] have shown that extremely light PBHs could usher in an era of early matter-domination (eMD) and source a large gravitational wave (GW) background when they evaporate, during the rapid transition from matter-domination to radiation-domination (RD). Although generally undetectable by present-day GW experiments, these GWs would constitute an excess contribution to the radiation energy density. As a result, $\Delta N_{\rm eff}$ limits derived from CMB and BBN observations constrain the abundance of PBHs down to $M\gtrsim 10^{-2}\textrm{g}$ [42, 43].
@@
In this paper, we consider PBH reformation during and after the eMD era. Shortly after matter-radiation equality, the overdensities associated with the PBHs grow and could become substantial before they evaporate. During the eMD phase, the probability of PBH reformation is enhanced, following a power law [46, 47, 48, 49, 50] instead of the usual exponential dependence [2, 51]. This enables the possibility that collections of PBHs may collapse and form much larger PBHs. A similar concept was explored in Ref. [52], where they discussed mergers of solar mass PBHs into supermassive black hole seeds, enhanced by strong clustering. In contrast, we focus on very light PBHs that would otherwise evaporate, and could initiate an eMD era in the early Universe, with collapse of a PBH gas rather than individual mergers as the formation mechanism. These reformed PBHs would be much lighter, actively emitting Hawking radiation which could be detected along with the coincident GW signal. This reformation framework is similar to PBH formation from other compact objects, such as Q-balls [53, 54] and oscillons [21, 55], with sufficient initial density to initiate an eMD phase and decay rate for the initial population to evaporate.
@@
We start by reviewing the relevant quantities and scales for PBH formation and the onset of an eMD era in Sec. II. The power spectrum and its growth is discussed in Sec. III which leads to the PBH reformation during eMD in Sec. IV. The accompanying GW background is discussed in Sec. V. We then discuss the additional possibility of reformation during the RD phase after PBH evaporation in Sec. VI. Finally, we conclude in Sec. VII.
@@ II PBH Parameters
$M_{\rm PBH}=\gamma\frac{4\pi M_{P}^{2}}{H}\simeq 47~{}\textrm{g}~{}\left(\frac{\gamma}{0.5}\right)\left(\frac{T}{10^{15}\textrm{ GeV}}\right)^{-2}~{},$ (1)
where we take $\gamma=0.5$ as a fiducial value [56] for both the initial formation and the reformation. Unless otherwise specified, this relation is used for the following calculations, but should be modified if the initial formation scenario involves subhorizon scales (e.g. from bubble collisions [7, 8, 10] or compact remnants [11, 12, 14, 13]).
At their formation, the fraction of PBHs with respect to the overall energy density is characterized by $\beta=\rho_{\rm PBH}/\rho_{\rm tot}$, where $\rho_{\rm tot}$ is the total energy density of the PBHs and the relativistic SM plasma. Here we differentiate between the density fraction of PBHs at initial formation, $\beta_{\rm if}$, and the density fraction at reformation $\beta_{\rm rf}$. We focus on initial formation masses, $M_{\rm pl}<M_{\rm if}\lesssim 4\times 10^{8}\textrm{g}$, which are unconstrained by BBN and CMB considerations. Such light PBHs are typical of GUT-scale formation [57, 58], bounce cosmology [59, 60], and preheating scenarios [61, 62, 63, 64].
Since PBHs are non-relativistic objects, their density increases relative to the SM plasma with redshift, which can result in an eMD phase. The matter(PBH)-radiation equality scale factor is simply related to the scale factor at initial formation by $a_{\rm eq}=a_{\rm if}/\beta_{\rm if}$. Likewise, the Hubble rate at the equality is proportional to the Hubble rate at initial formation, and the comoving horizon wavenumber is given by
$k_{\rm eq}=a_{\rm eq}H_{\rm eq}=\sqrt{2}a_{\rm if}H_{\rm if}\beta_{\rm if}~{}.$ (2)
@@
$\tau_{\rm evap}=4.0\times 10^{-4}\textrm{s}\left(\frac{M_{\rm PBH}}{10^{8}\textrm{ g}}\right)^{3}\left(\frac{108}{g_{H}}\right)~{},$ (3)
where the Hawking radiation degrees of freedom $g_{H}\simeq 108$ incorporates all the SM modes for $M_{\rm PBH}\lesssim 10^{13}\textrm{g}$.
@@
$\beta_{\rm if}>4\times 10^{-6}\left(\frac{M_{\rm PBH}}{1\textrm{ g}}\right)^{-1}\left(\frac{g_{*}(T_{\rm RH})}{g_{*}(T_{\rm if})}\right)^{1/3}~{},$ (4)
@@
×$\times\left(\frac{M_{\rm PBH}}{1~{}\textrm{g}}\right)^{-3/2}\left(\frac{g_{*}(T_{\rm RH})}{106.75}\right)^{-1/4}\left(\frac{g_{H}}{108}\right).$
@@
$=$ $2\times 10^{-4}\,a_{\rm if}H_{\rm if}\left(\frac{\beta}{\gamma}\right)^{1/3}\left(\frac{M_{\rm PBH}}{1~{}\textrm{g}}\right)^{-2/3}.$
@@ III Power Spectrum
The mean separation between PBHs at the initial formation, $\bar{r}$, defines the comoving ultraviolet cut-off wavenumber [41, 42],
$k_{\rm UV}=\frac{a}{\bar{r}}=\left(\frac{\beta_{\rm if}}{\gamma}\right)^{1/3}a_{\rm if}H_{\rm if}~{}.$ (7)
For $k<k_{\rm UV}$, the PBHs can be treated as a Poissonian fluid with initial density power spectrum [41]
$\mathcal{P}_{\delta}(t_{\rm if})=\frac{2}{3\pi}\left(\frac{k}{k_{\rm UV}}\right)^{3}~{}.$ (8)
During the eMD, PBH density perturbations grow linearly with scale factor [65] once they cross the horizon. On the other hand, the growth of short wavelength modes that enter during RD is delayed until matter-radiation equality.
The PBH density perturbation starts as an isocurvature mode at the moment of initial formation. But it becomes the adiabatic perturbation once PBHs dominate the Universe, similar to the curvaton scenario [66, 67, 68, 69]. Therefore, PBH reformation from the blue-tilted power spectrum in Eq. (8) is comparable to PBH formation in curvaton models with substantial small-scale power [70, 71].
The evolution of the power spectrum is described by the transfer function, $\mathcal{P}_{\delta}(t)=\mathcal{T}^{2}(t)\mathcal{P}_{\delta}(t_{\rm if})$. We numerically computed the linear transfer function under the relevant initial conditions (see Supplemental Material), and found a semi-analytic fit deep in the eMD phase,
@@
$\simeq\begin{cases}\frac{3}{2}\,y,&\text{for $\kappa\gg 1$}\\ \frac{4}{15}\,\kappa^{2}y,&\text{for $\kappa\ll 1$}\end{cases}~{}.$
Here $\kappa=k/k_{\rm eq}$ and $y=a/a_{\rm eq}$ are the wavenumber and scale factor with respect to the matter-radiation equality, and $y\gg 1$ is assumed.
@@
$\sigma^{2}_{\delta}(r,t)=\int^{\infty}_{0}\mathcal{P}_{\delta}(t)W^{2}(kr)\frac{dk}{k}~{}.$ (10)
@@
$W^{2}(kr\gg 1)\simeq\frac{9}{2}\frac{1}{(kr)^{4}}\simeq\frac{9}{8}\frac{1}{\kappa^{4}y^{2}}~{}$ (12)
@@
$P_{\delta}(t)W^{2}(kr)\propto\begin{cases}\kappa^{-1},&\text{for $\kappa>1$}\\ \kappa^{3},&\text{for $\kappa<1$}\end{cases}~{}$ (13)
with time dependence canceled for a long enough period of eMD.
Figure 1: The initial PBH power spectrum (gray), power spectrum at the moment of reheating (black) and at $y=y_{\rm rh}/10$ (black dashed), and the variance at the horizon scale per logarithmic interval of wavenumber (red). The initial power spectrum is given by the initial Poisson distribution of PBHs, which then grows according to the transfer function, and then multiplied by the window function to give the variance. The parameters are chosen to produce a sizeable population of reformed PBHs, represented by Case A in Table 1 and Fig. 2. All four lines are truncated at the UV cutoff $k_{\rm UV}$ in Eq. (7).
In Fig. 1, we show the power spectrum at the initial formation time and during the eMD era based on linear analysis. As in Eq. (13), the variance per logarithmic wavenumber intervals is constant in time and levels off at the matter-radiation equality scale $k\simeq k_{\rm eq}$. If any initial clustering of PBHs exists, the power at small scales increases and the contribution to the variance continues to increase for larger wavenumbers, making the behavior around the ultraviolet cutoff dominate the reformation.
At large $k$ deep in the matter-dominated era, the linear power spectrum at subhorizon scales can reach $P_{\delta}\gtrsim\mathcal{O}(1)$ which marks the breakdown of the linear theory. Just like particle dark matter, PBHs will typically form halos at those scales (see Ref. [72] for another reformation scenario through mergers in these halos), undergoing non-linear collapse and increasing in density by an $\mathcal{O}(100)$ factor [73]. We neglect these non-linear effects, which may increase the overdensity at horizon scales, and instead use linear theory to obtain a conservative estimate of the PBH reformation probability. Although there are well-known methods [74, 75, 76, 77, 78, 79, 80] to analytically transform the linear power spectrum into the non-linear power spectrum, we find that they do not yield reasonable results in this application. We leave non-linear studies to future works.
@@ IV PBH Reformation in eMD
Refs. [46, 47, 48, 49, 50] considered the possibility of PBH formation during MD. In the absence of pressure forces, the critical overdensity threshold may be expected to tend toward zero [81]. However, two additional considerations of isotropy and homogeneity are required to ensure gravitational collapse. Instead of forming a PBH, an anisotropic perturbation undergoes a “pancake” collapse, forming caustics and resulting in a virialized configuration. Using the anisotropy distribution of Ref. [82], the fraction of perturbations that are sufficiently spherically symmetric enough to collapse into a PBH was computed by Ref. [46, 47, 48], well approximated by $\beta=0.05556\sigma_{\delta}^{5}$. These exceptional horizon-scale perturbations avoid virialization and halo formation.
Additionally, these perturbations must be homogeneous enough to form a horizon (i.e. satisfy the hoop criterion) before forming a “naked singularity” which signifies the breakdown of the calculation. This limitation on the structure is satisfied by only a fraction of perturbations $\beta=3.70\sigma_{\delta}^{3/2}$ [50], but may be model dependent [48]. Taken independently, these two conditions would suggest a minimum PBH formation probability proportional to $\beta\propto\sigma^{13/2}$. However, both conditions are more easily satisfied by larger perturbations, which can shrink to their larger Schwarzchild radii before becoming too anisotropic or inhomogenous. Since these conditions are correlated and dependent, the actual conditional probability should then be much higher than the minimum estimate. Therefore, we use the optimistic estimate of $\beta\simeq 0.05556\sigma^{5}$ in this paper, which assumes that perturbations satisfying the much more stringent constraint of isotropy are also homogeneous enough. We ignore spin and angular momentum effects [49] which may alter the probability of PBH formation for small values of $\sigma_{\delta}$.
@@
Figure 2: Two populations of PBHs reformed during the eMD era that could be detected in the next generation of experiments. Both Case A (red) and Case B (blue) are in the allowed region of initial PBH parameters (green region) where the initial density $\beta_{\rm if}$ is large enough to result in an eMD phase and below the GW bound. The PBH spectrum for Case A lies below the current bounds from BBN and the emission of PBHs reformed in Case B could be detected by the next generation of CMB and $\gamma$-ray experiments. The gray-shaded regions are constraints from BBN [28], CMB [38, 39], extragalactic $\gamma$-rays [28], galactic $\gamma$-rays [29], Voyager-1 $e^{\pm}$ [30], and AMS-02 [31].
We apply the formula of Ref. [48] to PBH reformation during the eMD, from matter-radiation equality up until evaporation and reheating. Evaluating the variance at the horizon size of each epoch by Eqs. (10) and (13), we find it to be roughly constant and approximated by
@@
$f_{\rm PBH}=4.0\times 10^{20}\beta\left(\frac{M_{\rm if}}{1~{}\textrm{g}}\right)^{-3/2}\left(\frac{g_{*}(T_{\rm RH})}{106.75}\right)^{1/12}~{}.$ (15)
This flat mass spectrum has endpoints defined by the horizon scales at matter-radiation equality and reheating, $k_{\rm RH}<k<k_{\rm eq}$. From Eqs. (1), (2), and (5), the mass range of the reformed PBH spectrum is
$\frac{M_{\rm if}}{\sqrt{2}\beta_{\rm if}^{2}}\leq M_{\rm RH}\leq 6.0\times 10^{10}\,\textrm{g}\left(\frac{\gamma}{0.5}\right)\left(\frac{M_{\rm if}}{1~{}\textrm{g}}\right)^{3}~{}.$ (16)
@@
For Case A, the reformed PBH population are at the cusp of current BBN constraints from observations of the primordial deuterium abundance [83, 84, 85]. Experiments targeting the processes relevant for deuterium production could further reduce prediction uncertainties and lead to either more stringent constraints or an anomaly [86, 87, 88]. The reformed PBH population of Case B have lifetimes comparable to that of the Universe $\sim 10^{10}$ yr, and could be emitting high energy particles ($\gamma$,$e^{\pm}$) at the present. The evaporation signal of these reformed PBH may be within the sensitivity limit of the operational run of LHAASO [34, 89] and next generation $\gamma$-ray experiments such as CTA [35] and SGSO [36]. We consider a possible coincident GW signature in the next section.